### v1

# Experiential Semantic Information and Brain Alignment: Are Multimodal Models Better than Language Models?

Abstract
A common assumption in Computational Linguistics is that text representations learnt by multimodal models are richer and more human-like than those by language-only models, as they are grounded in images or audio—similar to how human language is grounded in real-world experiences. However, empirical studies checking whether this is true are largely lacking. We address this gap by comparing word representations from contrastive multimodal models vs. language only ones in the extent to which they capture experiential information—as defined by an existing norm-based ‘experiential model’—and align with human fMRI responses. Our results indicate that, surprisingly, language-only models are superior to multimodal ones in both respects. Additionally, they learn more unique brain-relevant semantic information beyond that shared with the experiential model. Overall, our study highlights the need to develop computational models that better integrate the complementary semantic information provided by multimodal data sources.
Experiential Semantic Information and Brain Alignment:
Are Multimodal Models Better than Language Models?
Anna Bavaresco, Raquel Fernández Institute for Logic, Language and Computation University of Amsterdam {a.bavaresco, raquel.fernandez}@uva.nl
1 Introduction
The relationship between abstract linguistic representations and the real-world entities they refer to is central to the academic discourse around semantics—the ‘symbol-grounding problem’ Harnad (1990); Bender and Koller (2020). While some researchers think that word meanings should be intended as purely symbolic Fodor (1983), others believe that words have meanings precisely because they are linked to specific entities, experiences or notions Barsalou (2008). Supporters of the latter view stress that human language acquisition is situated in a rich multimodal environment, where new words are learnt through interactions with objects and people Vigliocco et al. (2014). Theories of embodied cognition further highlight the importance of linking words to concrete experience not only for their acquisition but also for their comprehension. Indeed, according to these theories, understanding sentences involves engaging perceptual, motor or emotional simulations of their content (for an overview, see Kaschak et al., 2024).
The idea of obtaining richer semantic representations by learning them from sources other than text, such as images or audio, has inspired a great deal of computational work, from early attempts at concatenating image and text embeddings (e.g., Bruni et al., 2014; Kiela and Bottou, 2014; Derby et al., 2018; Davis et al., 2019) to the most recent large vision-language models (LVLMs, e.g., Li et al., 2023; Wang et al., 2024; Liu et al., 2024; Deitke et al., 2024; Laurençon et al., 2024). Some of these works aimed to obtain language representations aligning more closely with human responses, such as similarity judgments, while others were more oriented towards improving performance on benchmarks or downstream applications. Regardless of the end goal, all these works present multimodality as a desideratum, assuming that images provide additional semantic information that cannot be learnt from text alone; however, there is little to no work investigating which these semantic aspects are. In this paper, we aim to fill this gap by addressing the following question: Do multimodal models learn some facets of meaning related to perceptual experiences that language-only models cannot capture?
Operationalising the ‘extra-linguistic’ information that multimodal models are allegedly learning is a prerequisite for approaching this issue. We did this by relying on a semantic model introduced by Fernandino et al. (2022) to capture ‘experiential information’. This cognitive model represents words as $n$-dimensional arrays where each entry corresponds to aggregated human ratings on a pre-defined experiential attribute (e.g., Vision, Motion, Harm). We focused on a set of nouns and evaluated the alignment between their representations provided by the experiential model and those by comparable unimodal (language-only) and multimodal (vision-language and audio-language) computational models. This analysis allowed us to uncover if multimodal models indeed reflect more semantic information than language-only models. Next, we checked whether capturing experiential information translates into higher alignment with fMRI brain responses to the same set of noun words.
Our findings indicate several interesting trends. First, both vision-language and language-only models exhibit significant alignment with the experiential model and brain responses, while the audio-language model displays weak or non-significant correlations. Second, this alignment is more pronounced for language-only models, which appear to capture a great deal of brain-relevant information beyond experiential. Lastly, language-only models remain more brain-aligned than vision-language models even when focusing on a set of more concrete words, although the gap is reduced. Overall, our study shows that current multimodal models learn less brain-relevant information—both experiential and beyond—than comparable language-only models, highlighting the need to explore different approaches to construct multimodal word representations. The code to reproduce our experiments will be released upon publication.
2 Background
2.1 Embodied cognition
Embodied cognition identifies a suite of theoretical frameworks holding that language is understood by perceptual, emotional, or motor simulations of its content (e.g., Barsalou, 1999; Glenberg and Gallese, 2012; Zwaan, 2014; Pulvermüller, 2018). This general principle has received empirical support from multiple studies, both behavioural and neuroscientific.
For example, a series of works on the Action-sentence Compatibility Effect (ACE, Glenberg and Kaschak, 2002) and its subsequent variants Borreggine and Kaschak (2006); Zwaan and Taylor (2006); Bub and Masson (2012) revealed a significant difference in reaction times—attributed to motor simulations—when participants had to respond to a sentence (e.g., You passed the note to Art) with a movement (extending their arm) matching vs.  non-matching (retreating their arm) that mentioned in the sentence. Similarly, the sentence-picture verification task Stanfield and Zwaan (2001), where participants have to respond to a picture that is either compatible (an eagle with its wings outstretched) or incompatible (an eagle with its wings folded) with a sentence (The eagle is in the sky), and its variations Connell (2007); Hoeben Mannaert et al. (2017) have also been widely used to demonstrate the occurrence of perceptual simulation during language comprehension. In parallel, a line of neuroscientific studies have found evidence that semantic processing may activate motor (among others, Hauk et al., 2004; Tettamanti et al., 2005; Aziz-Zadeh et al., 2006) and perceptual brain regions Kiefer et al. (2008); Van Dam et al. (2012).
2.2 Multimodal models of semantics
Embodied cognition and related ideas, such as visual grounding, have percolated from Cognitive Science to Computational Linguistics, motivating attempts to build semantic models that learn representations from data sources beyond text. Early efforts in this direction (e.g., Bruni et al., 2014; Kiela and Bottou, 2014; Lazaridou et al., 2015; Silberer and Lapata, 2012, 2014) were characterised by 1) a focus on developing human-aligned computational models of meaning and 2) limited computational modelling resources available (large datasets of paired image-text inputs did not exist at the time, nor did large transformer-based architectures).
Recently, multimodal models have become more powerful and found application on a variety of downstream tasks (e.g., image captioning, image retrieval, or visual question answering). Some seminal works used a contrastive objective to learn aligned image and text representations Radford et al. (2021); Jia et al. (2021), while others—often inspired by BERT’s Devlin et al. (2019) successes in language modelling—tried to apply its underlying intuitions to the vision-language domain Tan and Bansal (2019); Li et al. (2019); Lu et al. (2019); Chen et al. (2020). Finally, state-of-the-art large vision-language models (Li et al., 2023; Wang et al., 2024; Liu et al., 2024; Deitke et al., 2024; Laurençon et al., 2024, LVLMs, e.g.,), which usually bridge together a large language model (LLM) and an image encoder, can engage in strikingly human-like conversations about images. In contrast to the early attempts at multimodal modelling, these works share 1) a focus on solving, or improving performance on, downstream tasks, and 2) the availability of massive datasets and large models with billions of parameters.
For our experiments, we aimed to leverage models that are powerful while, at the same time, suitable for drawing cognitively-meaningful comparisons. The need to satisfy both constraints prevented us from evaluating state-of-the-art LVLMs; we elaborate more on our model choices in Section 3.2.
2.3 Experiential models of semantics
Recently, a few approaches motivated by embodied cognition have introduced models of semantics aimed at capturing ‘experiential information’, i.e., aspects of meaning related to how humans ground language in experiences. These experiential models were constructed by asking human annotators to rank words on a set of pre-defined dimensions. For example, Fernandino et al. (2022) introduced an experiential model based on 48 dimensions spanning perceptual, emotional, and action-related constructs. In two fMRI studies, they found that the experiential model yielded more brain-aligned word representations than taxonomic and distributional models; additionally, it contributed unique semantic information not represented by the other models.
Similarly, Carota et al. (2024) experimented with a different experiential model based on 11 dimensions and compared its brain alignment against that of a distributional model. Their study revealed significant correlations with brain responses in more ROIs (regions of interest) for the experiential model than for the distributional model. However, an integrative model combining both displayed significant correlations in an even larger number of ROIs, suggesting that experiential and distributional are complementary aspects of human semantic processing.
Despite their merits, experiential models are bounded in their accuracy by an a priori selection of dimensions and, relying on human annotations, remain expensive to construct. These limitations open the intriguing question of whether experiential information can be captured by computational models learning semantic representations in a data-driven fashion.
3 Methods
3.1 Data and experiential model
For our experiments, we used word stimuli, fMRI responses and experiential model from Study 2 by Fernandino et al. (2022). [^1] These materials have been made publicly available by Fernandino et al. The full list of words and the experiential features can be found at https://www.pnas.org/doi/10.1073/pnas.2108091119#supplementary-materials; fMRI data are available at https://osf.io/87chb/. We describe each below.
Word stimuli
Word stimuli comprise 320 nouns, half of which refer to objects and the other half to events. The 160 object nouns include an equal number of words (40) from four categories (food, vehicles, animals, tools); likewise, the event nouns span four semantic subcategories (social event, negative event, sound, communication).
fMRI responses
fMRI responses were collected from $36$ participants while they viewed the above-mentioned word stimuli one at a time and rated the frequency with which they experienced their corresponding entities in daily life. We focused on voxels from a ‘semantic network ROI’ defined by Binder et al. (2009) based on a meta-analysis and averaged responses across participants.
Experiential model
The experiential model, hereafter abbreviated as Exp48, represents each word as a set of ratings on 48 pre-defined dimensions capturing different aspects of people’s experience with objects/events, e.g., Vision, Hand action or Unpleasant. The ratings were introduced by Binder et al. (2016) as part of a wider set of experiential salience norms; they range from 0 to 6 and were provided by 1743 unique crowdworkers.
3.2 Computational models
Our model choices were motivated by the goal to maximise comparability across architectures. More concretely, we selected three models (language-only, vision-language, and audio-language) comparable in terms of fine-tuning objective—the contrastive one—and architecture—they all have a pretrained BERT Devlin et al. (2019) as language encoder. [^2] All three models were released with both BERT-based Devlin et al. (2019) and RoBERTa-based Liu et al. (2019) implementations. We used the former in all our experiments. One aspect in which these architectures differ is the amount of training data; however, we believe this issue does not invalidate our results and further discuss it in Limitations (Appendix A).
SimCSE
(Simple Contrastive Learning of Sentence Embeddings, Gao et al., 2021) is a language-only sentence encoder fine-tuned contrastively on 1M sentences randomly sampled from English Wikipedia. Matching pairs for the contrastive objective were created by applying different dropout masks to the same sentence.
MCSE
(Multimodal Contrastive Learning of Sentence Embeddings, Zhang et al., 2022) is a vision-language sentence encoder fine-tuned by jointly optimising a SimCSE objective and a CLIP-like Radford et al. (2021) objective. The fine-tuning data for the first objective was the same as SimCSE’s; as for the CLIP-like objective, where a matching pair was defined by an image and its caption, the fine-tuning data consisted of 83K images from MS-COCO Lin et al. (2014) annotated with multiple captions.
CLAP
(Contrastive Language Audio Pretraining, Wu et al., 2023) is an audio-language model whose language encoder was initialised with pre-trained BERT weights and fine-tuned on audio-caption pairs with a CLIP-like objective. The fine-tuning data included $633,526$ audio-text pairs, with audio clips representing human activities, natural sounds, and audio effects.
For reference, we also tested BERT and VisualBERT Li et al. (2019) as its visual counterpart.
BERT
Devlin et al. (2019) is a transformer-based language-only model pretrained with two objectives: masked language modelling and next sentence prediction. Its pretraining data included the BooksCorpus (800M words, Zhu et al., 2015) and English Wikipedia (2500 words). As mentioned above, SimCSE, MCSE and CLAP fine-tuned pretrained BERT architectures.
VisualBERT
Li et al. (2019) is a vision-language model consisting of a BERT-based language encoder (initialised with parameters from pretrained BERT) and a pretrained visual feature extractor based on Faster RCNN (Ren et al., 2015). Its training objectives, which mirror BERT’s, were masked language modelling with image input and sentence-image prediction. The vision-language pretraining data comprised MS-COCO and VQA 2.0 Goyal et al. (2017). Note that this is not a contrastive model; we included it for reference as it can be considered as a vision-language extension of BERT, but it is not directly comparable with MCSE, SimCSE and CLAP.
Refer to caption Refer to caption
Figure 1: Results from representational similarity analysis. On the left, Spearman correlations quantifying the alignment between word representations from Exp48 and by computational models. On the right, Spearman correlations indicating the alignment between fMRI responses from human participants and word representations by computational models.
3.3 Extracting representations
Given that all the models we considered were trained to learn contextualised representations from sentences, single words may be an out-of-distribution input. Therefore, following an approach similar to May et al. (2019), we embedded the noun stimuli in a set of generic template sentences (e.g., Someone mentioned the [word], see Appendix B for the complete list) when passing them to the models. [^3] We empirically verified that passing words within templates rather than in isolation yields higher alignment with both the experiential model and brain responses (see Appendix C.1). For all templates, we derived word representations from the hidden states of each layer; more specifically, we selected the hidden states corresponding to the tokens of the target word and averaged them across templates.
3.4 Alignment evaluation
To compare model representations against Exp48 and brain responses, we used representational similarity analysis (RSA, Kriegeskorte et al., 2008), which quantifies the alignment between two representational spaces (either by two models or by a model and brain responses) as the correlation between representational dissimilarity matrices (RDMs). In our experiments, RDMs were populated with pairwise cosine distances between model representations or fMRI responses for all the unique word-pairs. The alignment between RDMs was calculated as a Spearman correlation.
RSA allows comparing models’ alignment with Exp48 or brain responses, but it does not reveal whether models explain shared variance or provide independent contributions. Fernandino et al. (2022) computed partial correlations to check how much brain-relevant information Exp48 shared with the other models they considered, i.e., two distributional models (Word2vec and GloVe; Mikolov et al., 2013; Pennington et al., 2014) and two taxonomic models (a WordNet-based model and a categorical one). We used the same approach to determine how much brain-relevant information our tested models share with Exp48 and with each other. Partial correlations were obtained by regressing the RDM from one model out of the RDM from another model and then computing a Spearman correlation between the residuals and the RDM from brain responses.
4 Results
4.1 Exp48 and brain alignment across models
We performed RSA to obtain a first measure of model representations’ alignment with Exp48 and fMRI responses. This analysis was conducted on model representations averaged across the three layers yielding the highest alignment individually; note that these layers may differ when considering alignment to brain responses vs. Exp48 (see Appendix C.2 for a visualisation of layer-wise alignment). The results from RSA against brain responses and Exp48 are displayed in Figure 1. All Spearman correlations are statistically significant ($p$ < 0.05), except for CLAP’s correlation with brain responses ($p$ = 0.70); we additionally verified that all the pairwise differences between correlations are statistically significant. [^4] Statistical significance was determined by applying a Fisher transformation to the correlation coefficients from each pair of models and calculating the $p$-value associated with the difference between the two $z$-scores. All $p$-values were Bonferroni-corrected with $\alpha$ = 0.05. The same approach for verifying statistical significance was applied to all correlation comparisons throughout the paper.
An inspection of correlations against Exp48 indicates BERT as the most aligned model ($\rho$ = 0.53); SimCSE and MCSE also display moderate correlations with Exp48 ($\rho$ = 0.52 and $\rho$ = 0.45, respectively). In contrast, CLAP’s representations are poorly aligned with Exp48, exhibiting a correlation of just 0.03. A comparison between vision-language models (MCSE and VisualBERT) and their unimodal counterparts (SimCSE and BERT) reveals that the former, surprisingly, reflect less experiential information than the latter.
Regarding alignment with brain responses in the semantic ROI, BERT is again the best model ($\rho$ = 0.23), although it remains less brain-aligned than Exp48 ($\rho$ = 0.27). All the other models display positive correlations, with the exception of CLAP, whose correlation is not statistically significant ($\rho$ = 0.00, $p$ = 0.70). Similarly to the Exp48-alignment results, here we found that the language-only models BERT and SimCSE are more brain-aligned than their vision-language extensions VisualBERT and MCSE. We delve deeper into the robustness of this finding in Section 5.
Refer to caption Refer to caption
Figure 2: Results from partial correlation analyses. On the left, Spearman correlations between brain responses and the residuals obtained regressing model RDMs out of the Exp48 RDM. The dotted line indicates Exp48’s initial brain alignment without removing any information. On the right, Spearman correlations between brain responses and the residuals obtained regressing the Exp48 RDM out of model RDMs. The bars in lighter shades indicate models’ initial brain alignment.
An interesting trend common across results from both RSAs (against Exp48 and fMRI responses) is that representations by SimCSE and MCSE—which have been shown to outperform BERT on semantic text similarity tasks Gao et al. (2021); Zhang et al. (2022)—are less aligned than those by BERT. A potential explanation for this may be that we considered single-word representations. Since contrastive fine-tuning, as applied to SimCSE and MCSE, optimises sentence-level representations as opposed to token-level ones, it could be that some token-level semantic properties initially learnt by BERT got somehow diluted through this process.
4.2 Experiential information vs. unique contribution in models’ brain alignment
Results from the partial correlation analysis are displayed in Figure 2, whose left-hand panel shows how much Exp48 representations align with brain responses without the information they share with each of the other models. An interesting observation is that the lowest correlations were obtained when regressing out BERT and SimCSE. This provides an interesting complement to the findings from RSA against Exp48 representations: RSA shows that BERT and SimCSE share substantial representational information with Exp48, and partial correlations suggest that this information is also brain-relevant. Regarding models’ brain alignment without Exp48, displayed in Figure 2’s right-hand panel, a noteworthy finding is that BERT’s and SimCSE’s representations are the most brain-aligned even after regressing out Exp48. This suggests that these models learnt some semantic information that is not captured by Exp48 but is still reflected in brain responses.
Additionally, for each model we checked which proportion of its initial brain alignment is attributable to unique contribution as opposed to information shared with Exp48. This can be visualised by comparing the dark-shade bars against the light-shade ones in the right-hand panel of Figure 2. An interesting result revealed by this comparison is that, although MCSE is more brain-aligned than VisualBERT, their unique contribution without Exp48 is the same in absolute value ($\rho$ = 0.06); in other terms, 50% of VisualBERT’s brain alignment is due to unique information, while in MCSE it is 32%. Regarding BERT and SimCSE, the majority of their initial brain alignment is eroded when regressing out Exp48; however, the asymmetry is not substantial, and the unique contribution accounts for more than 40% of the initial brain alignment in both models. As for CLAP, it exhibits a weak negative correlation that is not statistically significant, confirming that the model does not contribute any brain-relevant information.
Finally, we used partial correlations to compare vision-language models (VLMs) against their language-only counterparts (LMs). We found that neither MCSE ($\rho$ = 0.00; $p$ = 0.60) nor VisualBERT ($\rho$ = 0.00; $p$ = 0.66) exhibit statistically significant correlations with brain responses once SimCSE and BERT, respectively, are regressed out. Crucially, this indicates that VLMs did not learn any additional brain-relevant information besides that already captured by their LM counterparts.
5 Assessing Results’ Robustness
RSA results revealed a consistent advantage of language-only models over the multimodal ones. This finding contrasts with the expectation—shared across a great deal of work on multimodality and language modelling—that training models on diverse data modalities, as opposed to text alone, should yield more human-like language representations. In the following, we present two analyses aimed at assessing the robustness of these findings. Given that the audio-language model CLAP did not achieve a statistically significant brain alignment, we excluded it from further analyses and focused on the remaining vision-language and language-only architectures.
5.1 Do caption-like templates result in improved brain alignment?
As pointed out by Tan and Bansal (2020), image captions are examples of grounded language, which differs from other types of natural language along many dimensions. Since the VLMs we evaluated were trained on image-caption pairs, they may have over-fitted to the language present in captions. Therefore, it is possible that the sentence templates we used to obtain contextualised word representations from the models are somehow out-of-distribution for VLMs.
To control for this potential confound, we re-extracted word representations employing different templates whose structure was modelled around captions (e.g., There is an [object] in a [place], or A [person] is [verb in -ing] in a [place]). These structures were identified based on a manual inspection of captions from MS-COCO, which was part of both MCSE’s and VisualBERT’s training. Given the challenges of creating caption-like templates providing a fitting context for all the word stimuli, we used different sets of templates for each sub-category of words described in Section 3.1 (e.g., There is a [food-word] on a table in a restaurant or A few people gathered for a [social event-word]). We provide the complete list of templates in Appendix B.
The procedure for calculating brain alignment was the same as that employed in the main experiment. Spearman correlations between model-derived RDMs and the fMRI-derived RDM are displayed in Figure 3. All correlations are statistically significant, as well as correlation differences between models. A comparison across models confirms the trend from the main experiment: Language-only models are more brain-aligned than their vision-language counterparts. This suggests that the finding is robust and not a by-product of the templates where word stimuli were embedded.
The dotted lines in Figure 3 allow comparing the brain alignment model representations achieve when using caption-like templates vs. when using the templates from the main experiment. This comparison reveals that all models—not only VLMs—exhibit higher brain alignment when using caption-like templates. We interpret this as indicating that caption-like templates are not more in-distribution for VLMs, but rather provide a better-specified context that is beneficial to all models.
Figure 3: Spearman correlations indicating alignment between model representations extracted using caption-like templates and fMRI responses. Dotted lines indicate the initial correlations obtained with the templates from the main experiment.
5.2 Do VLMs yield more brain-aligned representations for objects vs. events?
Provided that VLMs learn additional semantic information, it could be that not all word representations benefit from multimodal training to the same extent; instead, a potential advantage may be more prominent for words referring to visual contents. The words used in our main experiments include nouns from multiple semantic categories (see Section 3.1 for more details), which may largely vary in their degree of ‘visual-ness’. Therefore, it is possible that we did not detect additional brain-relevant information learnt by VLMs because we focused on the ‘wrong’ words.
To check whether this is the case, we further analysed two word subsets with different levels of concreteness. The subsets were identified by leveraging the semantic labels already present in our word set, i.e., objects and events. [^5] In their supplementary materials, Fernandino et al. (2022) report that the average concreteness score for objects is 4.9, while for events it is 3.6. We repeated RSA separately for these two word subsets following the same procedure employed in the main experiment.
The results of this analysis are displayed in Figure 4. A first observation is that—for all models except VisualBERT—correlations are statistically significantly stronger for events than objects. This pattern was also reported by Fernandino et al. (2022), who attributed it to “higher variability of pairwise similarities for the neural representations of event concepts”.
A second interesting result is that the model ranking we observed analysing the entire word set (BERT > SimCSE > MCSE > VisualBERT) is replicated for events but not for objects, where none of the differences between model correlations is statistically significant. While there is a negative effect overall, further training BERT on image-text pairs (as in VisualBERT) or fine-tuning it with a contrastive objective (as in SimCSE and MCSE) does not significantly alter the initial brain alignment of its object-word representations. Interestingly, Exp48, which we included for reference, is outperformed by BERT on events; however, it remains statistically significantly more brain-aligned than the other models on objects.
Finally, comparing vision-language models against their language-only counterparts shows that BERT and VisualBERT do not significantly differ regarding the brain alignment of their object-word representations, while SimCSE and MCSE do (with SimCSE remaining more aligned). [^6] Note that, since we used Bonferroni corrections, this difference is statistically significant here—but not when comparing all five models—due to a change in the number of relevant comparisons (2 vs. 5). As for event-word representations, SimCSE and BERT are, respectively, significantly more brain-aligned than MCSE and VisualBERT. These results further support the robustness of our initial finding, i.e., that LMs models are more aligned than their VL counterparts. However, the reduced gap between the two model types when considering object-word representations vs. event-word ones suggests that VLMs do, comparatively, learn more brain-aligned representations for objects than events.
Figure 4: Spearman correlations indicating alignment between model representations and fMRI responses. Correlations are displayed separately for object-words and event-related words.
6 Discussion and Conclusions
Our study provides an in-depth comparison between multimodal and language-only architectures in terms of their ability to capture experiential semantic information and their alignment with brain responses. While multimodal models are often expected to learn additional semantic aspects that language-only models cannot learn, our results reveal that their word representations are less aligned with Exp48 and fMRI responses than those by LMs. Moreover, within multimodal models, the vision-language ones show moderate positive correlations with Exp48 and fMRI responses, while the audio-language one correlates weakly with Exp48 and does not yield a significant correlation with brain responses. A potential explanation for these counterintuitive findings is that the dimensions used to create Exp48 are moderately abstract, whereas the extra information learnt by multimodal models may concern lower-level features or patterns of co-occurrence. Alternatively, it could be that multimodal training is not as effective during fine-tuning as it would be in the pre-training stage. However, we verified that not even the language encoder from the powerful CLIP Radford et al. (2021)—pretrained contrastively on 400M image-text pairs and excluded from the main experiment as it is not directly comparable with the other architectures—yields more brain-aligned word representations than BERT and SimCSE (we report these results in Appendix C.3).
Another noteworthy result is that the correlations with fMRI responses we found for SimCSE, MCSE and BERT are higher than those achieved by the computational models (GloVe and Word2vec) tested in Fernandino et al. (2022) (see Appendix C.3). This finding aligns with previous work showing that transformer-based architectures are more predictive of brain responses during language processing than word-level embedding models and recurrent neural networks Schrimpf et al. (2021).
Finally, one last interesting finding was that the LMs and, to a larger extent, the VLMs we tested learn brain-relevant semantic information beyond that captured by Exp48. This partially echoes the results by Carota et al. (2024), with the difference that the computational model included in their study was strictly distributional.
Our study has several implications for future work. First, it invites caution against assuming that technical innovations allowing models to solve additional downstream tasks should necessarily make them more ‘human-like’. Second, it indicates that there is significant room for improving current computational language models so that they learn the brain-relevant experiential information they currently lack—how to concretely achieve this remains an open question.
Appendix
Appendix A Limitations
Despite our efforts to develop a robust experimental design, a few limitations should still be considered when drawing conclusions from our results. First, the experiential model we used—while showed to reflect brain-relevant information—relies on pre-defined dimensions that are, to some extent, abstract and high-level. In this sense, it may fail to capture low-level perceptual information relevant for modelling human word representations and, perhaps, learnt by computational models.
Another limitation is that our study focuses on single words as opposed to longer text passages. To some extent, our experiments suggest that this may affect machine language processing, as we found representations by computational models to be more aligned with brain responses when words are passed within sentence templates vs. in isolation (see also Appendix C.1). Similarly, the isolated-word setup is likely to play a role human language processing—as insightfully observed by Zwaan (2014), context is crucial in determining the mental simulations people engage during language comprehension and the amount of perceptual detail they contain.
Lastly, two additional limitations are related to the models included in our study. The first is that, while SimCSE, MCSE and CLAP are comparable in their learning objectives and architecture, they still differ in the amount of fine-tuning data. While, in principle, this could be problematic, we observed that SimCSE—despite being fine-tuned on less sentences than MCSE—still proves to be more Exp48- and brain-aligned. As for CLAP, we acknowledge that the smaller amount of fine-tuning audio-caption pairs, together with the fact that it did not optimise a SimCSE-like objective jointly with the CLIP-like one, may have played a role in its poor performance.
The second limitation related to our model selection is that we only considered architectures fine-tuned contrastively, which are known to reflect high-level correspondences between objects present in images and captions, while failing to capture more fine-grained relations between modalities Hendricks and Nematzadeh (2021); Parcalabescu et al. (2022); Thrush et al. (2022); Liu et al. (2023); Chen et al. (2023); Bavaresco et al. (2024). Therefore, it could be that different trends would emerge if repeating the experiments with computational models trained with different objectives.
Appendix B Sentence Templates
The neutral sentence templates where the word stimuli were embedded in order to obtain contextualised representations from the computational models were the following:
Someone mentioned the <word>.
The post was about the <word>.
Everyone was talking about the <word>.
They were all interested in the <word>.
People know about the <word>.
In one of our additional experiments (see Section 5.1), we used caption-like sentences to check whether they were more in-distribution for vision-language models and, therefore, yielded more Exp48- and brain-aligned representations. Below, we report the caption-like templates used for each word sub-category.
Templates used for the sub-category food:
There is a <word> on a table in a restaurant.
A <word> is on a kitchen table.
A woman is eating a <word>.
A <word> with a few glasses around.
A close-up of a <word>.
Templates used for the sub-category vehicle:
There is one man in a <word>.
A <word> is surrounded by a few people.
A woman is posing next to a <word>.
A <word> with a young man next to it.
A close-up of a <word>.
Templates used for the sub-category tool:
There is a man holding a <word>.
A <word> is lying on the ground.
A woman is using a <word>.
A <word> with some people in the background.
A close-up of a <word>.
Templates used for the sub-category animal:
There is a <word> eating voraciously.
A man is feeding a <word>.
A woman next to a <word>.
A <word> with a little girl staring at it.
A close-up of a <word>.
Templates used for the sub-category negative event:
There is a crowd looking scared because of a <word>.
Many people are trying to shelter from a <word>.
A <word> happening in a big city.
A <word> with many people involved.
A picture of a <word>.
Templates used for the sub-category social event:
There is a small crowd attending a <word>.
A few people are gathered for a <word>.
A <word> attended by a large group of people.
A <word> with many people involved.
A picture of a <word>.
Templates used for the sub-category communication:
There is a small crowd at a <word>.
A few people are participating in a <word>.
A <word> in a crowded room.
A <word> with many people involved.
A picture of a <word>.
Templates used for the sub-category sound:
There is a man hearing a <word>.
A few people seem to hear a <word>.
A <word> is heard by a few people.
A <word> with a few people listening to it.
A picture of a <word>.
Appendix C Additional RSA Results
Refer to caption Refer to caption
Figure 5: Spearman correlations observed from model representations obtained by passing single words vs. words embedded in templates. The left-hand panel shows the alignment with Exp48 and the right-hand one with brain responses.
C.1 Single-word vs. contextualised representations
Our choice to derive word representations by including them in sentences was guided by the intuition that single words could have been an out-of-distribution input for computational models trained to output contextualised word representations. We empirically verified that representations obtained by embedding words within templates yield higher alignment than those obtained by passing single words to the models. We show the Exp48 and brain alignment obtained with both embedding-extraction procedures in Figure 5.
C.2 Layer-wise RSA results
In the main paper, we reported RSA results calculated from model representations averaged across the three layers yielding the highest alignment individually. Here, we provide a layer-wise visualisation of RSA results, which allows observing how Exp48 vs.  brain alignment changes throughout model layers. Specifically, layer-wise Spearman correlations against Exp48 are displayed in Figure 6, while those against fMRI responses are in Figure 7.
Figure 6: Spearman correlations indicating how representational similarity between model representations and Exp48 representations changes along model layers.
Figure 7: Spearman correlations indicating how representational similarity between model representations and brain responses changes along model layers.
C.3 RSA with additional baselines
For completeness, in Table 1 we report RSA results including three additional models: CLIP Radford et al. (2021), a vision-language model pretrained contrastively on 400M image-caption pairs, and the distributional models GloVE Pennington et al. (2014) and Word2vec Mikolov et al. (2013). The distributional models were originally included in Fernandino et al. (2022); note that the brain correlations we report differ from the ones from Fernandino et al. (2022), as they computed an average across participant-wise brain correlations, while we averaged brain RDMs across participants before computing correlations.
Model $\rho$ Exp48 $\rho$ Brain
SimCSE 0.52 0.22
MCSE 0.45 0.19
CLAP 0.03 0.00
BERT 0.53 0.23
VisualBERT 0.27 0.12
CLIP 0.41 0.14
GloVe 0.45 0.14
Word2vec 0.42 0.125
Table 1: Spearman correlations quantifying the alignment of models’ representational spaces with Exp48 and brain responses.